# xDeepServe: Model-as-a-Service on Huawei CloudMatrix384

### v5

@@ 3.2 All-to-All Communication for Colocated MoE-Attention
Figure 7: Pull-based Dispatch based on UB’s Global Shared Memory. Dashed red lines are metadata transfer. Bulk data transfer in step 6 and 7 can also use the DMA engine. A single dispatch kernel can use multiple computing cores. All NPUs involved in a dispatch run the same seven steps described above. We show three NPUs for simplicity.
@@
This section presents the design and implementation of two all-to-all communication primitives—dispatch and combine—used in large-scale expert parallelism (EP). In EP, dispatch routes each token’s hidden state to its top-$k$ experts based on gating scores, while combine aggregates expert outputs, weighted by the same scores, into a unified tensor. Together, we find these primitives account for at least 25% of MoE execution time. Similar to point-to-point primitives, we implement these all-to-all operations using far memory semantics rather than traditional network-level verbs like DeepEP deepep . For brevity, we detail the dispatch protocol below; combine follows a similar pattern.
@@
Data Structure. Similar to the point-to-point primitives, we partition each NPU’s on-chip memory into three regions. The metadata area contains 32-byte fields, one per rank. Each field includes an event ID for sanity checking, a pointer to the corresponding rank’s buffer offset in the managed data area, and a token count received from that rank. The max number of metadata fields is similar to that of point-to-point primitives. The managed data area is partitioned by rank ID, with each NPU assigned a fixed-size block determined by the maximum batch size.
@@
Distributed Memory Protocol of dispatch. The dispatch protocol consists of two main phases: broadcasting the number of tokens each rank should receive, followed by the actual data transfer. Figure 7 illustrates the complete workflow. Step 1: The sender’s serving engine invokes dispatch, passing the source buffer in the app data area, an event ID, and additional parameters. A kernel is launched on the sender NPU, which uses mtu-in to copy data from the app data area to each computing core’s data buffer in parallel. Step 2: If quantization is enabled, the kernel converts the data from FP16/BF16 to INT8 using certain instructions. Step 3: The kernel writes token data into the managed data area, partitioned by destination rank ID. Step 4: The kernel updates each destination rank’s metadata with the number of tokens it will receive. Step 5: Each NPU polls its local metadata area until metadata from all ranks is received. Steps 6–7: Once all metadata is received, the kernel pulls token data from peer NPUs into the data buffer using mtu-in, guided by received offsets and token counts. It then copies the data into the destination buffer in the app data area via mtu-out.
@@
Performance. Figure 6 evaluates dispatch and combine’s performance under varying batch sizes per die, using DeepSeek-R1 models with a fixed EP128 configuration. At small batch sizes, dispatch exhibits slightly higher latency than combine due to the additional overhead from quantization. However, since quantization reduces data size by half, dispatch becomes faster than combine when the batch size per die exceeds 32. For reference, with a batch size per die of 96 and EP128, the total global batch size across the system reaches 96 × 128 = 12,288.
@@ 3.3 All-to-All Communication for Disaggregated MoE-Attention
This section presents the design of two all-to-all communication primitives—attention2expert (A2E) and expert2attention (E2A)—used in disaggregated MoE and attention deployments (see §5.2). These primitives enable efficient token routing and aggregation between attention NPUs and expert NPUs at scale.
@@
A key difference, however, is that attention and expert modules reside on separate NPU dies, introducing asymmetry in resource allocation. In large-scale deployments, this asymmetry presents new challenges. For example, in a typical DeepSeek-R1/V3 liu2024deepseek-v3 configuration with 288 experts, we often provision 288 expert NPUs but only 160 attention NPUs (see §5.2). In a naive pull-based dispatch design, each attention NPU would push metadata for all expert NPUs and wait for them to pull data. This approach quickly becomes inefficient due to the high fan-out and limited scalar throughput of each computing core.
@@
Trade-off between Memory Semantic and DMA. To improve communication efficiency, we employ NPU-Direct Unified Remote Memory Access (URMA), a technique on Ascend NPUs similar to IBGDA on GPUs liu2024deepseek-v3 . NPU-Direct URMA enables computing cores to issue remote memory access requests directly to the DMA engine, bypassing both the host CPU and AI CPU as shown in Figure 2.2. Although NPU-Direct URMA incurs higher startup latency compared to memoty transfer units, it offers three key advantages. First, it reduces computing core consumption. By offloading memory access to the DMA engine using asynchronous semantics, more computing core resources become available for communication or computation. Second, it is better suited for high-throughput scenarios. Memory transfer units transfers are limited by the computing core’s data buffer size, whereas the DMA engine supports transfers up to several GBs. Third, it avoids contention with compute streams. Since mtu-in is also used for computation, using DMA helps prevent interference when compute and communication streams share the same NPU die, as discussed in §5.2.
Performance. We evaluate A2E and E2A at SuperPod scale using a single deployment with three DP domains, each containing 160 DP groups (TP = 1), and 288 expert NPUs. We run DeepSeek-R1 with a per-die batch size of 96, resulting in a global batch size of $96\times 3\times 160=46{,}080$. Under this configuration, A2E achieves a latency of 172 µs, while E2A completes in 193 µs. These results demonstrate that both primitives maintain low-latency execution even under high concurrency and large-scale deployment.
@@ 4 Scalable Serving System at SuperPod-Scale
Figure 9: The Architecture of FlowServe. A single FlowServe engine can span an entire CloudMatrix384 SuperPod—48 Ascend NPU servers with 768 NPU dies. FlowServe scales at the granularity of a DP group, where each group encapsulates a complete serving pipeline. To avoid bottlenecks and single points of failure, both request scheduling and response handling are fully distributed across DP groups.
@@ 4.1 Overview
To this end, we redesign our serving engine, FlowServe hu2025deepflow , evolving it from a single-node deployment to a SuperPod-scale distributed system, as illustrated in Figure 9 and Figure 10. This redesign centers on three key components:
  • First, we introduce the Data Parallel (DP) group abstraction, inspired by SGLang sglang . Each DP group encapsulates a full serving pipeline—including tokenization, API parsing, SPMD-style executors, the Relational Tensor Cache (RTC), and the DistFlow networking stack hu2025deepflow . To eliminate bottlenecks and single points of failure, both request scheduling and response handling are fully decentralized across DP groups.
@@ 4.3 DP Load Balancing
Decode DP Load Balancing. In the decode phase, each DP group’s load is shaped by two factors: the number of concurrent requests it handles and the memory consumed by the KV cache. Each DP group supports a fixed batch size. When full, it must block incoming requests until others finish, which increases both time-to-second-token (TTST) and TPOT. At the same time, KV cache usage reflects the memory footprint of active sequences. Uneven usage across groups leads to unbalancd MLA execution time as shown in Figure 10 or even trigger swapping, severely degrading performance. To balance decode load, we first exclude DP groups that have reached their batch limit. Among the rest, we select the group with the lowest KV cache usage, accounting for reserved space needed for long outputs. The TE-shell tracks real-time metrics for each group: it updates the pending request count on dispatch and completion, and collects periodic KV cache stats. This design enables informed, system-wide scheduling decisions without introducing significant overhead.
@@ 4.4 Proactive Garbage Collection to Reduce Jitter
(a) We show the expert load distribution of a DeepSeek-R1 layer under the ShareGPT workload. The distribution is highly skewed—20% of experts receive more than the average load, and the hottest expert sees 30× more tokens than the average.
(b) The setup uses EP288 and 1K sequence. MoE-Avg-Routing, which forces uniform load across all experts; MoE-Native, which uses the original token-to-expert assignment; and MoE-Balanced, which applies our EPLB to balance expert load.
@@ 4.5 Expert Load Balancing
To address expert load imbalance, we use a data-driven approach. Our design periodically analyzes token routing patterns, identifies overloaded ("hot") experts, and replicates these experts across multiple NPUs. Each MoE layer reserves redundant slots per NPU to host replicated experts. During inference, we evenly distribute tokens to these replicas using precomputed mappings. Meanwhile, we continuously collect and analyze activation data to update expert replica assignments. Expert weights are swapped asynchronously to maintain uninterrupted inference throughput under significant load imbalance.
@@
Figure 12 illustrates how logical experts are mapped to rotated physical replicas based on token positions. Consider a case with 4 tokens per inference step and 256 logical experts. The mapping table has shape (4, 256), where each row corresponds to a token and each column to a logical expert. Suppose Logical Expert 1 has two replicas: a primary in slot 2 and a redundant in slot 1. The first column of the mapping table rotates between slots 1 and 2 across the 4 tokens, assigning each with equal probability to ensure balanced routing.
@@ 4.6 Multi-Token Prediction
This section describes how FlowServe integrates MTP into its disaggregated prefill-decode workflow (§5) to improve decode efficiency. Figure 13 illustrates the full execution pipeline. DeepSeek’s Multi-Token Prediction (MTP) liu2024deepseek-v3 extends conventional next-token prediction by training the model to generate multiple future tokens in sequence. During inference, MTP enables speculative decoding by predicting several tokens in one pass and verifying them afterward, accelerating autoregressive generation.
Execution Workflow. In the prefill stage, both the main model and MTP process the input prompt to construct the key-value (KV) cache and generate the first token. The KV cache and relevant hidden states are then transferred to the decode stage. During decoding, we run a tightly optimized five-step loop designed to eliminate CPU-induced latency bubbles. Our initial implementation followed the EAGLE framework from vLLM vllm-sosp23 , but its default scheduling introduced noticeable stalls. We replaced it with a customized pipeline that maximizes NPU utilization and minimizes idle time. The loop proceeds as follows: (1) Run MTP forward to generate k draft tokens; (2) Sample token candidates from the MTP outputs; (3) Verify the draft tokens using the main model; (4) Sample again from the main model outputs; (5) Check final logits to decide token acceptance.
@@ 4.7 INT8 Quantization of DeepSeek Models
The Ascend NPU in CloudMatrix384 does not natively support FP8 arithmetic. To deploy DeepSeek-R1/V3—originally trained in FP8—we quantize the model to INT8 using Post-Training Quantization (PTQ) method. Our method integrates SmoothQuant smoothquant-icml23 and GPTQ frantar2023gptq techniques specifically for DeepSeek models.
@@
MLA Quantization. The MLA module uses low-rank compression for the Q, K, and V matrices and integrates RoPE. We quantize important weights to INT8, including query compression (Wq_a), key/value compression (Wkv_a), query reconstruction (Wq_b), and attention output projection (Wo). Our analysis shows activations have a significantly wider dynamic range (10-100×) compared to weights. To handle this, we apply a smoothing operation that redistributes quantization difficulty between activations and weights before quantization. Additionally, we use a GPTQ-based channel-wise quantization method with Hessian-guided iterative refinement. This approach dynamically updates the remaining FP weights to compensate for quantization errors. Figure 15 shows the input activation and weight magnitude in a DeepSeek-R1 linear layer, before and after smoothing to reduce quantization outliers.
@@ 5 Transformerless: Towards Fully Disaggregated LLM Serving
We advocate resource disaggregation as a core architectural principle for modern model-serving systems. Disaggregation separates system components into independently scalable units, improves fault isolation, and enables flexible evolution of system software. Network bandwidth is the foundation of disaggregation. A 40 Gb/s scale-out network once enabled memory farm-nsdi14 and OS disaggregation legoos-osdi18 for traditional workloads. Today, 400 GB/s scale-up interconnects make it possible to disaggregate model serving for AI workloads.
To realize these benefits, we introduce Transformerless, an architecture that decomposes transformer-based LLMs into modular blocks—attention, feedforward (FFN), and MoE layers—executed independently on NPUs interconnected via high-speed fabric. The hundreds of GB/s UB interconnect in CloudMatrix384 enables aggressive disaggregation without sacrificing inference performance. We envision Transformerless in three stages. First, in disaggregated prefill-decode (§5.1), we isolate compute-bound prefill from memory-bound decode by assigning them to different NPUs. Second, in disaggregated MoE-attention (§5.2), we decouple MoE experts from attention computation, enabling independent scaling by sequence length and batch size. Finally, in dataflow serving at SuperPod scale (§5.3), we aim to break global synchronization altogether, allowing truly independent scaling of each serving component.
5.1 Disaggregated Prefill and Decode
Figure 16: The Workflow of Disaggregated Prefill and Decode over CloudMatrix384. We support $M$ prefill and $N$ decode deployments with full-mesh connectivity. We illustrate the end-to-end workflow of sending a request from a prefill TE to a decode TE.
Disaggregated prefill and decode were concurrently introduced by Splitwise patel2023splitwise , TetriServe tetriserve-arxiv24 , and DistServe zhong2024distserve to reduce contention between the compute-bound prefill and memory-bound decode phases. This design was further refined by Mooncake qin2024mooncake and MemServe hu2024memserve .
Despite its adoption, building an efficient and robust disaggregated prefill/decode pipeline remains challenging—especially for large MoE models with expert parallelism—due to two main reasons. First, prefill and decode require distinct parallelization strategies because of differing sequence lengths: we use TP=4 for prefill attention and TP=1 for decode. This necessitates different DP groupings, making colocated execution inefficient. Second, the backend execution diverges: prefill uses an eager graph for dynamic-length support, sequence parallelism, and prefix cache computation; decode uses a static graph for maximum throughput.

### v6

@@ 3.2 All-to-All Communication for Colocated MoE-Attention
Figure 7: Pull-based Dispatch based on UB’s Global Shared Memory. Dashed red lines are metadata transfer. Bulk data transfer in step 6 and 7 can also use the DMA engine. A single dispatch kernel can use multiple AIV cores. All NPUs involved in a dispatch run the same seven steps described above. We show three NPUs for simplicity.
@@
This section presents the design and implementation of two all-to-all communication primitives—dispatch and combine—used in large-scale expert parallelism (EP). In EP, dispatch routes each token’s hidden state to its top-$k$ experts based on gating scores, while combine aggregates expert outputs, weighted by the same scores, into a unified tensor. Together, we find these primitives account for at least 25% of MoE execution time. Similar to point-to-point primitives, we implement these all-to-all operations using far memory semantics rather than traditional network-level verbs like DeepEP [deepep]. For brevity, we detail the dispatch protocol below; combine follows a similar pattern.
@@
Data Structure. Similar to the point-to-point primitives, we partition each NPU’s on-chip memory into three regions. The metadata area contains 32-byte fields, one per rank. Each field includes an event ID for sanity checking, a pointer to the corresponding rank’s buffer offset in the managed data area, and a token count received from that rank. The max number of metadata fields is similar to that of point-to-point primitives. The managed data area is partitioned by rank ID, with each NPU assigned a fixed-size block determined by the maximum supported batch size.
@@
Distributed Memory Protocol of dispatch. The dispatch protocol consists of two main phases: broadcasting the number of tokens each rank should receive, followed by the actual data transfer. Figure 7 illustrates the complete workflow. Step 1: The sender’s serving engine invokes dispatch, passing the source buffer in the app data area, an event ID, and additional parameters. A kernel is launched on the sender NPU, which uses MTE2 to copy data from the app data area to each AIV’s unified buffer in parallel. Step 2: If quantization is enabled, the kernel converts the data from FP16/BF16 to INT8 using vector instructions. Step 3: The kernel writes token data into the managed data area, partitioned by destination rank ID. Step 4: The kernel updates each destination rank’s metadata with the number of tokens it will receive. Step 5: Each NPU polls its local metadata area until metadata from all ranks is received. Steps 6–7: Once all metadata is received, the kernel pulls token data from peer NPUs into the unified buffer using MTE2, guided by received offsets and token counts. It then copies the data into the destination buffer in the app data area via MTE3.
@@
Performance. Figure 6 evaluates dispatch and combine’s performance under varying batch sizes per die, using DeepSeek-R1 models with a fixed EP128 configuration. At small batch sizes, dispatch exhibits slightly higher latency than combine due to the additional overhead from quantization. However, since quantization reduces data size by half, dispatch becomes faster than combine when the batch size per die exceeds 32. For reference, with a batch size per die of 96 and EP128, the total global batch size across the system reaches $96\times 128=12{,}288$.
@@ 3.3 All-to-All Communication for Disaggregated MoE-Attention
This section presents the design of two all-to-all communication primitives—A2E (attention2expert) and E2A (expert2attention)—used in disaggregated MoE-Attention deployments (see §5.2). These primitives enable efficient token routing and aggregation between attention NPUs and expert NPUs at scale.
@@
A key difference, however, is that attention and expert modules reside on separate NPU dies, introducing asymmetry in resource allocation. In large-scale deployments, this asymmetry presents new challenges. For example, in a typical DeepSeek-R1/V3 [liu2024deepseek-v3] configuration with 288 experts, we often provision 288 expert NPUs but only 160 attention NPUs (see §5.2). In a naive pull-based dispatch design, each attention NPU would push metadata for all expert NPUs and wait for them to pull data. This approach quickly becomes inefficient due to the high fan-out and limited scalar throughput of each AIV core.
@@
Trade-off between MTE and DMA. To improve communication efficiency, we employ NPU-Direct Unified Remote Memory Access (URMA), a technique on Ascend NPUs similar to IBGDA on GPUs [liu2024deepseek-v3]. NPU-Direct URMA enables AIV cores to issue remote memory access requests directly to the DMA engine, bypassing both the host CPU and AI CPU as shown in §2.2. Although NPU-Direct URMA incurs higher startup latency compared to MTE2/MTE3, it offers three key advantages. First, it reduces AIV consumption. By offloading memory access to the DMA engine using asynchronous semantics, more AIV resources become available for communication or computation. Second, it is better suited for high-throughput scenarios. MTE2/MTE3 transfers are limited by the AIV’s unified buffer size, whereas the DMA engine supports transfers up to several GBs. Third, it avoids contention with compute streams. Since MTE2 is also used for computation, using DMA helps prevent interference when compute and communication streams share the same NPU die, as discussed in §5.2.
Performance. We evaluate A2E and E2A at SuperPod scale using a single deployment with three DP domains, each containing 160 DP groups (TP = 1), and 288 expert NPUs. We run DeepSeek-R1 with a per-die batch size of 96, resulting in a global batch size of $96\times 3\times 160=46{,}080$. Under this configuration, A2E achieves a latency of 172 $\mu s$, while E2A completes in 193 $\mu s$. These results demonstrate that both primitives maintain low-latency execution even under high concurrency and large-scale deployment.
@@ 4 Scalable Serving System at SuperPod-scale
Figure 9: The Architecture of FlowServe. A single FlowServe engine can span an entire CloudMatrix384 SuperPod—48 Ascend 910C servers with 768 NPU dies. FlowServe scales at the granularity of a DP group, where each group encapsulates a complete serving pipeline. To avoid bottlenecks and single points of failure, both request scheduling and response handling are fully distributed across DP groups.
@@ 4.1 Overview
To this end, we redesign our serving engine, FlowServe [hu2025deepflow], evolving it from a single-node deployment to a SuperPod-scale distributed system, as illustrated in Figure 9 and Figure 10. This redesign centers on three key components:
  • First, we introduce the Data Parallel (DP) group abstraction, inspired by SGLang [sglang]. Each DP group encapsulates a full serving pipeline—including tokenization, API parsing, SPMD-style executors, the Relational Tensor Cache (RTC), and the DistFlow networking stack [hu2025deepflow]. To eliminate bottlenecks and single points of failure, both request scheduling and response handling are fully decentralized across DP groups.
@@ 4.3 DP Load Balancing
Decode DP Load Balancing. In the decode phase, each DP group’s load is shaped by two factors: the number of concurrent requests it handles and the memory consumed by the KV cache. Each DP group supports a fixed batch size. When full, it must block incoming requests until others finish, which increases both time-to-second-token (TTST) and TPOT. At the same time, KV cache usage reflects the memory footprint of active sequences. Uneven usage across groups leads to unbalanced MLA execution time as shown in Figure 10 or even triggers swapping, severely degrading performance. To balance decode load, we first exclude DP groups that have reached their batch limit. Among the rest, we select the group with the lowest KV cache usage, accounting for reserved space needed for long outputs. The TE-shell tracks real-time metrics for each group: it updates the pending request count on dispatch and completion, and collects periodic KV cache stats. This design enables informed, system-wide scheduling decisions without introducing significant overhead.
@@ 4.4 Proactive Garbage Collection to Reduce Jitter
[…]
(b) The setup uses EP288 and 1K-token sequence length. MoE-Avg-Routing, which forces uniform load across all experts; MoE-Native, which uses the original token-to-expert assignment; and MoE-Balanced, which applies our EPLB to balance expert load.
@@ 4.5 Expert Load Balancing
To address expert load imbalance, we use a data-driven approach. Our design periodically analyzes token routing patterns, identifies overloaded (“hot”) experts, and replicates these experts across multiple NPUs. Each MoE layer reserves redundant slots per NPU to host replicated experts. During inference, we evenly distribute tokens to these replicas using precomputed mappings. Meanwhile, we continuously collect and analyze activation data to update expert replica assignments. Expert weights are swapped asynchronously to maintain uninterrupted inference throughput under significant load imbalance.
@@
Figure 12 illustrates how logical experts are mapped to rotated physical replicas based on token positions. Consider a case with 4 tokens per inference step and 256 logical experts. The mapping table has shape [4, 256], where each row corresponds to a token and each column to a logical expert. Suppose Logical Expert 1 has two replicas: a primary in slot 2 and a redundant in slot 1. The first column of the mapping table rotates between slots 1 and 2 across the 4 tokens, assigning each with equal probability to ensure balanced routing.
@@ 4.6 Multi-Token Prediction
DeepSeek’s Multi-Token Prediction (MTP) [liu2024deepseek-v3] extends conventional next-token prediction by training the model to generate multiple future tokens in sequence. During inference, MTP enables speculative decoding by predicting several tokens in one pass and verifying them afterward, accelerating autoregressive generation.
This section describes how FlowServe integrates MTP into its disaggregated Prefill-Decode workflow (§5) to improve decode efficiency. Figure 13 illustrates the full execution pipeline.
Execution Workflow. In the prefill stage, both the main model and MTP process the input prompt to construct the key-value (KV) cache and generate the first token. The KV cache and relevant hidden states are then transferred to the decode stage. During decoding, we run a tightly optimized five-step loop designed to eliminate CPU-induced latency bubbles. Our initial implementation followed the EAGLE framework from vLLM [vllm-sosp23], but its default scheduling introduced noticeable stalls. We replaced it with a customized pipeline that maximizes NPU utilization and minimizes idle time. The loop proceeds as follows: (1) Run MTP forward to generate k draft tokens; (2) Sample token candidates from the MTP outputs; (3) Verify the draft tokens using the main model; (4) Sample again from the main model outputs; (5) Check final logits to decide token acceptance.
@@ 4.7 INT8 Quantization of DeepSeek Models
The Ascend 910C NPU in CloudMatrix384 does not natively support FP8 arithmetic. To deploy DeepSeek-R1/V3—originally trained in FP8—we quantize the model to INT8 using a Post-Training Quantization (PTQ) method. Our method integrates SmoothQuant [smoothquant-icml23] and GPTQ [frantar2023gptq] techniques specifically for DeepSeek models.
@@
[…]
@@
Figure 16: The Trend Towards Transformerless Serving. Our system evolves from a PD-colocated setting to disaggregated Prefill-Decode, then to disaggregated MoE-Attention, and ultimately into a fully asynchronous dataflow serving system.
@@ 5 Transformerless: Towards Fully Disaggregated LLM Serving
We advocate resource disaggregation as a core architectural principle for modern model-serving systems. Disaggregation separates system components into independently scalable units, improves fault isolation, and enables flexible evolution of system software. Network bandwidth is the foundation of disaggregation. A 40 Gb/s scale-out network once enabled memory [farm-nsdi14] and OS disaggregation [legoos-osdi18] for traditional workloads. Today, 400 GB/s scale-up interconnects make it possible to disaggregate model serving for AI workloads.
To realize these benefits, we introduce Transformerless, an architecture that decomposes transformer-based LLMs into modular blocks—attention, feedforward (FFN), and MoE layers—executed independently on NPUs interconnected via high-speed fabric. The hundreds of GB/s UB interconnect in CloudMatrix384 enables aggressive disaggregation without sacrificing inference performance. We envision Transformerless in three stages (Figure 16). First, in disaggregated Prefill-Decode (§5.1), we isolate compute-bound prefill from memory-bound decode by assigning them to different NPUs. Second, in disaggregated MoE-Attention (§5.2), we decouple MoE experts from attention computation, enabling independent scaling by sequence length and batch size. Finally, in dataflow serving at SuperPod scale (§5.3), we aim to break global synchronization altogether, allowing truly independent scaling of each serving component.
5.1 Disaggregated Prefill-Decode
Figure 17: The Workflow of Disaggregated Prefill-Decode over CloudMatrix384. We support $M$ prefill and $N$ decode deployments with full-mesh connectivity. We illustrate the end-to-end workflow of sending a request from a prefill TE to a decode TE.
Disaggregated Prefill-Decode was concurrently introduced by Splitwise [patel2023splitwise], TetriServe [tetriserve-arxiv24], and DistServe [zhong2024distserve] to reduce contention between the compute-bound prefill and memory-bound decode phases. This design was further refined by Mooncake [qin2024mooncake] and MemServe [hu2024memserve].
Despite its adoption, building an efficient and robust disaggregated Prefill-Decode pipeline remains challenging—especially for large MoE models with expert parallelism—due to two main reasons. First, prefill and decode require distinct parallelization strategies because of differing sequence lengths: we use TP=4 for prefill attention and TP=1 for decode. This necessitates different DP groupings, making colocated execution inefficient. Second, the backend execution diverges: prefill uses an eager graph for dynamic-length support, sequence parallelism, and prefix cache computation; decode uses a static graph for maximum throughput.